Find the number of ways in which each commitee can be selected.
-3 people from a group of 5

Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(5,3) = (5!)/(3!*(5-3)!)

C(5,3) = (5!)/(3!*2!)

C(5,3) = (5*4*3!)/(3!*2!)

C(5,3) = (5*4)/(2!)

C(5,3) = (5*4)/(2*1)

C(5,3) = (20)/(2)

C(5,3) = 10


So there are 10 ways (assuming order doesn't matter)